# Synthesizing data products, mathematical models, and observational measurements for lake temperature forecasting

### v1

@@ Appendix A Toy problem
Simulated observations (gray dots), along with fits from all four models are shown in Figure 17, and 95% PIs and CIs are shown in Table 1. The predictive means from all four methods are nearly identical. We calculated CI coverage for methods 1-3 above in two ways. For moments-based methods 2 and 3 we calculated standard errors for CI/PI calculations using $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ from Eq. 12. For the hetGP implementation, our calculations were similar but involved dividing the diagonal of the predictive covariance matrix plus a varying nugget term by the number of replicates (15). We then determined the percentage of sample means that fell within CIs. To calculate PIs we did not divide by the number of replicates and report the percentage of data points that fell within the intervals. For this CI/PI calculation, performance is similar among the four methods with Vecchia B providing more conservative UQ by design. Vecchia B achieved 98% and 100% PI and CI coverage, respectively, while coverages of the other three methods ranged between 93 and 97% (Table 1, middle column).
@@
Here, $\Sigma_{\_}{ij}=k(q(\mathbf{x}_{\_}i,\mathbf{x}_{\_}j))+g\mathbb{I}_{\_}{\{i=j\}})$, $k(\cdot)$ is the Gaussian covariance kernel, and $\Sigma(\bar{\mathbf{X}}_{\_}n)$ is $n\times n$, where $\bar{\mathbf{X}}_{\_}n$ is the matrix of unique input locations as defined in Section 3.2. In Eq. (15), the quantity $\mu_{\_}n^{(v)^{*}}(\mathbf{x})$ is derived from the full GP fit to $D_{\_}n^{(v)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{S}}_{\_}n)$ via jmleGP from the laGP package. The quantity $\sigma_{\_}n^{2(m)}(\mathbf{x})$ either represents the predictive variance of the full GP fit to $D_{\_}n^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$ (the moments-based approach method 2) or predictive variance of the Vecchia-approximated fit to $D_{\_}n^{(m)}$ (moments-based approaches Vecchia A and B in method 3). While the SE in Eq. (15) is partially based on a full GP fit using a Gaussian covariance kernel, coverages for Vecchia A and B (which are fitted using a Matern kernel with a smoothness parameter of 3.5) calculated this way still represent good estimates for the purposes of this example. Calculating true confidence intervals with the heteroscedastic GP implementation in method 1) requires only the diagonal of the “no nugget” predictive variance, which is returned by mleHetGP.
@@
The final step involves an MCMC experiment in which CI coverages are calculated 100 times using Eqn. (15), and the proportion of times the true mean falls within the CI is reported. These coverage values are given in the right-most column of Table 1. All four methods perform well in this regard, with all achieving at least 97% coverage.

### v2

@@ A.2 Toy problem
Simulated observations (gray dots), along with fits from all four models are shown in Figure 12, and 95% PIs and CIs are shown in Table 1. The predictive means from all four methods are nearly identical. We calculated CI coverage for methods 1)–3) above in two ways. For moments-based methods 2) and 3) we calculated standard errors for CI/PI calculations using $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ from Eq. 11. For the hetGP implementation, our calculations were similar but involved dividing the diagonal of the predictive covariance matrix plus a varying nugget term by the number of replicates (15). We then determined the percentage of sample means that fell within CIs. To calculate PIs we did not divide by the number of replicates and report the percentage of data points that fell within the intervals. For this CI/PI calculation, performance is similar among the four methods with Vecchia B providing more conservative UQ by design. Vecchia B achieved 98% and 100% PI and CI coverage, respectively, while coverages of the other three methods ranged between 93 and 97% (Table 1, middle column).
@@
Here, $\Sigma_{\_}{ij}=k(q(\mathbf{x}_{\_}i,\mathbf{x}_{\_}j))+g\mathbb{I}_{\_}{\{i=j\}})$, $k(\cdot)$ is the Gaussian covariance kernel, and $\Sigma(\bar{\mathbf{X}}_{\_}n)$ is $n\times n$, where $\bar{\mathbf{X}}_{\_}n$ is the matrix of unique input locations as defined in Section 3.2. In Eq. (15), the quantity $\mu_{\_}n^{(v)^{*}}(\mathbf{x})$ is derived from the full GP fit to $D_{\_}n^{(v)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{S}}_{\_}n)$ via jmleGP from the laGP package. The quantity $\sigma_{\_}n^{2(m)}(\mathbf{x})$ either represents the predictive variance of the full GP fit to $D_{\_}n^{(m)}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}_{\_}n)$ (the moments-based approach method 2) or predictive variance of the Vecchia-approximated fit to $D_{\_}n^{(m)}$ (moments-based approaches Vecchia A and B in method 3). While the SE in Eq. (15) is partially based on a full GP fit using a Gaussian covariance kernel, coverages for Vecchia A and B (which are fitted using a Matérn kernel with a smoothness parameter of 3.5) calculated this way still represent good estimates for the purposes of this example. Calculating true confidence intervals with the heteroscedastic GP implementation in method 1) requires only the diagonal of the “no nugget” predictive variance, which is returned by mleHetGP.
@@
The final step involves a MC experiment in which CI coverages are calculated 100 times using Eq. (15), and the proportion of times the true mean falls within the CI is reported. These coverage values are given in the right-most column of Table 1. All four methods perform well in this regard, with all achieving at least 97% coverage.
A.3 Exercise with GLM runs
To benchmark our Vecchia-based workaround for SK we developed the following exercise to compare to true SK with respect to out of sample using a subset of NOAA-GLM runs. The goal was to have a large enough training exercise to stress our Vecchia calculations, while not being too large as to preclude expedient ordinary-GP calculations behind SK. We divided up one year’s worth of NOAA-GLM hindcasts, including all depths and horizons, into five-day chunks. For each chunk we used a random $n_{\_}i=16$ sample of NOAA-replicates for training, setting aside the other 15 for out-of-sample testing. In this way, each training data set was manageably sized at $n=24,000$ (10 depths $\times$ 30 horizons $\times$ 5 days $\times$ 16 ensemble members) unique inputs. For each of the seventy-three spans of five days in 2022 we fit SK (910) and the Vecchia-approximated first- and second-moments, $\mu_{\_}n^{(m)}$ and $\sigma^{2(m)}_{\_}n$ for the mean and $\mu^{(v)}_{\_}n$ for full predictive uncertainty in Eq. (11), respectively. We saved out-of-sample RMSEs and coverages of 95% CIs and PIs. The calculation of CI coverage involved averages of the fifteen held-out replicates, at each unique input, whereas PIs used all of the hold-out replicates individually.
Figure 13: Vecchia approximation (Vecchia SK) and true SK 95% confidence and prediction interval (CI and PI, respectively) coverages over day of year (DOY) in 2022.
Figure 13 provides a compact summary of the results of that experiment. PI/CI coverages are shown as a proportion. PIs for both the approximate and true SK are fairly consistent throughout the year, whereas CIs fluctuate slightly more over DOY. Coverages of both true SK and approximate SK vary between DOY 200 and 250. We attribute the slightly higher variability in CI coverage of both methods to the relatively small 5-day training data size that was entertained. With only five days of examples to train on, there is not much diversity along the time axis in order to stabilize variances of first moments. However, average CI coverages over the entire year for both methods are close to the nominal value of 95%. True SK and approximate SK achieved 97.2 and 93.8% coverage, respectively, while average PI coverages for true SK and approximate SK are 94.5 and 99%. Overall, coverage of our Vecchia-based SK approximation (11), via $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathbf{x})$ and $\hat{\sigma}^{\mathrm{SK}}(\mathbf{x})^{2}$ is slightly higher than that of SK, which is by design (11). Additionally, we found that the ratio of root mean-squared errors (RMSEs) from true SK and our Vecchia approximation, scaled by the range in temperatures at all depths and for all days in 2022, was near 0.035 for all days in 2022. This indicates that true SK and Vecchia are in agreement with respect to predictive accuracy.
Appendix B Iterative forecasting framework
Our forecasting is faithful to the brief description in Section 4.1, depicted diagrammaticaly in Figure 14, but is more accurately described by an operation that is performed all at once in a matter of minutes at midnight (00UTC), separating “yesterday” from “tomorrow”. This is when NOAA forecasts for “tomorrow” (and the 29 subsequent days), and sensor measurements from “yesterday” become available. Yesterday’s sensor measurements can be used to validate previous forecasts, and the metrics we prefer are discussed next in Section 5.1. NOAA forecasts can be fed into GLM to produce 1–30 day-ahead lake temperature forecasts over each horizon and depth of interest. Those NOAA-GLM forecasts are then incorporated into the corpus of existing simulation training data, along with the sensor measurements. Model fitting for surrogates and bias correction can then again be carried out in an updating step, for example as described in Sections 3.33.4. Although there is potential for computational economy in priming aspects of new fits from previous ones, we do not generally bother because the Vecchia-based methods are fast. Updated surrogates may then be used to generate a 1–30 day-ahead bias-corrected forecast. The process repeats again at the next midnight as part of the iterative, near-term forecasting cycle (Dietze et al., 2018).
Forecasting accuracy and needs vary over the year during different meteorological and hydrological conditions and in response to changing manager needs. More frequent forecasts may be necessary in the fall to anticipate mixing events that are associated with water quality degradation (Thomas et al., 2020). In comparison, when the lake water column is stratified in the summer and water temperatures are not rapidly changing, managers may require less frequent forecasts. Consequently, for our focal reservoir, it would be ideal if the updates to surrogates happened at a sub-daily time scale. However, in our effort to produce a holistic and robust forecasting apparatus, we opted for daily updates, which are generally sufficient for lake temperature forecasting.
Appendix C Scoring metrics
Given predictions $\hat{\mu}\equiv\hat{\mu}(\mathcal{X})=(\hat{\mu}_{\_}1,\dots,\hat{\mu}_{\_}N)^{\top}$ and predictive variance $\Sigma=\Sigma(\mathcal{X})$ for out of sample $Y\equiv Y(\mathcal{X})=(y_{\_}1,\dots,y_{\_}N)^{\top}$ score and RMSE may be calculated as follows:
For score, $\Sigma$ refers to the predictive variance of each method and is not fixed.
When validating for multiple horizons, and/or over multiple days, we aggregate accordingly. Score, RMSE, interval width and coverage are interesting because they provide a general sense of how the forecaster is performing. However, they are most useful when benchmarking against other, competing frameworks.
Appendix D Additional empirical results
Here we provide auxiliary empirical analysis for the forecasting experiment in Section 4.
D.1 Depth analysis
D.2 Breaking out three days of forecasts
Figure 16 augments the visual in Figure 11 by breaking out horizon and depth contributions to bias for three particular days (indicated as vertical dashed lines in Figure 11). These three example days were chosen to display occasions when the bias-correction was correct in sign but not in magnitude (DOY 48 and 260) and generally correct in sign and magnitude (171). Consider DOY 48 first, in the left panels of Figure 16.
Applying a bias correction (GPBC) improves the accuracy of forecasts compared to the GLM surrogate without a bias correction (GPGLM) with respect to observations (purple dots) at both 0 and 5m depths, although at 5m depth the forecast is still characterized by a cold bias. Moving to DOY 260 (right panel), the bias-correction actually worsens the forecast at the surface depth at after 15 days into the future. Finally, at day 171 (middle panel), forecasts at surface and 5m depths are generally improved by the bias-correction. These plots demonstrate that predicted bias applied to out-of-sample forecasts from GPBC mainly improves forecasts, although there are periods during the year when bias-correction over- or under-corrects or is of the wrong magnitude.
Appendix E Additional discussion
We suspect the main reason we lose to the climatological model (OGP) is a paucity of training data for estimating bias and the quality of NOAA forecasts used to drive GLM. We only have 2.5 years, meaning that we have seen fewer than three examples of yearly variation. Although OGP has access to the same amount of observational data, its simpler design – not requiring the correction of poor model-based forecasts at longer horizons – helps to prevent over-fitting. GPBC can be modified so it is less weather dependent and behaves more like a climatological model by omitting the input $\phi$ (GPBC without $\phi$). While it performs better at shorter horizons, GPBC with $\phi$ is still our preferred model, because overall, it still performs better than GPBC without $\phi$. Moreover, because the variability of GLM ensemble forecasts increases with forecast horizon, it is likely that the shape and degree of bias at longer horizons is more difficult to estimate compared to shorter horizons, where ensemble forecasts are more in agreement. Thus, bias correction at longer horizons may be less effective than at lower horizons without many more examples of yearly variation.
The quality of NOAA-GLM forecasts is largely dependent on the quality of NOAA forecasts themselves (assuming that GLM is well-calibrated). Weather forecast accuracy more than 10 days into the future declines quickly (Bauer, Thorpe and Brunet, 2015). Moreover, NOAA forecasted variables represent spatial averages over a 25km grid cell, much bigger than FCR for example, which may be too coarse to capture fine-scale variation in meteorological variables. Downscaling and bias-correction of NOAA forecasted variables could result in more accurate lake temperature forecasts from GLM, but additional post-processing of NOAA forecasts would take substantial computation time. Consequently, we decided that an investigation was beyond the scope of our current study. While forecasts of the climatological benchmark model were more accurate than GPBC at longer horizons, the UQ of those forecasts were generally poorer compared to those of GPBC.
It could be possible to create a hybrid forecasting model, in which the strength of $\phi(\lambda,t)$ is adjusted depending on the forecast horizon. Such an approach would combine the merits of GPBC with $\phi$ (better at shorter horizons) and GPBC without $\phi$ (more accurate at longer horizons). It is also worth noting that while OGP was only training on about 2.5 years of data, it performed exceedingly well at forecast horizons greater than 2 weeks. OGP shows that even very simple climatological models provide hard-to-beat benchmarks.
While the covariance for DOY should technically be periodic, we opted to use a standard (non-periodic) Gaussian covariance function for OGP. We investigated the use of a work-around to obtain a periodic fit, but we found that this did not result in any significant improvements to the non-periodic fit.
A way to extend this work would be to use GPBC as a likelihood for setting configuration parameters to GLM. In other words, to use the surrogate as a statistical calibration mechanism for the computer model (Kennedy and O’Hagan, 2001; Liu et al., 2009), as opposed to relying on expert opinion using observed meteorological data, not NOAA forecasts.
We could also treat the derived input $\phi(t,\lambda)$, allowing us to distinguish forecasts among years, (i.e., the weather inputs) as a latent variable that could be estimated along with other quantities, thus allowing us to distinguish forecasts among years. Such an approach, which involves estimating a high-dimensional free quantity for all $t$ and $\lambda$, is precluded, at the moment, by a paucity of sensor measurements. Perhaps once we have accumulated 10+ years of additional data , spanning more yearly environmental variability, the benefits of such an approach might outweigh the additional estimation risk it implies.
Another extension of this work could focus on extremes. Extreme events, such as heatwaves, rapid cooling events, or autumn mixing, are particularly critical for managing drinking water resources and mitigating water quality degradation. These events can lead to conditions such as hypoxia, algal blooms, or disruptions in water treatment processes. Accurately forecasting these extremes, as well as quantifying the uncertainty associated with them, is essential for operational decisions. For example, improved predictions during extreme mixing events could help managers adjust intake levels to minimize the impact on drinking water quality.
While our current methodology is designed for daily surrogate updates, incorporating extremes into the forecasting framework could involve future enhancements, such as prioritizing extremes in model training using weighted loss functions or cost-sensitive approaches, adopting adaptive surrogate models that increase temporal resolution (e.g., hourly) based on triggers such as rapid changes in meteorological conditions, exploring event-specific forecast evaluation metrics that better capture performance during rare but impactful scenarios.